OpenStudy (anonymous):

Which equation models the problem below if c represents the number of kittens whose eyes are closed? There are 22 kittens in a photograph, and 16 of them have their eyes open. How many kittens have their eyes closed?

OpenStudy (anonymous):

16+c=22 right?..because in total there's 22 kittens, nd 16 kittens eyes r open....so basically, the rest of the kittens eyes are close...so 22-16=6-(kittens eyes close)

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

ik..lol